# Bounding Boxes and Probabilistic Graphical Models: Video Anomaly Detection Simplified

### v1

@@ Abstract
In this study, we formulate the task of Video Anomaly Detection as a probabilistic analysis of object bounding boxes. We hypothesize that the representation of objects via their bounding boxes only, can be sufficient to successfully identify anomalous events in a scene. The implied value of this approach is increased object anonymization, faster model training and fewer computational resources. This can particularly benefit applications within video surveillance running on edge devices such as cameras. We design our model based on human reasoning which lends itself to explaining model output in human-understandable terms. Meanwhile, the slowest model trains within less than $7$ seconds on a $11^{th}$ Generation Intel Core i9 Processor. While our approach constitutes a drastic reduction of problem feature space in comparison with prior art, we show that this does not result in a reduction in performance: the results we report are highly competitive on the benchmark datasets CUHK Avenue and ShanghaiTech, and significantly exceed on the latest State-of-the-Art results on StreetScene, which has so far proven to be the most challenging VAD dataset.
@@ 1 Introduction
  • A new perspective on problem domain representation within VAD given the current solution landscape, and latest demands from industry.
  • A novel approach for fully object-centric VAD in which high-level object bounding box attributes are learned by a discrete Bayesian Network (BN) to detect visual outliers in video streams.
  • Human-understandable model and result explainability within the context of these high-level attributes used to model the problem domain.
@@ 3.1 Discrete Bayesian Network
In order to sufficiently represent the spatial dimension of the model which is responsible for localizing anomalous events within a single frame, we divide the image into a uniform grid structure of quadratic cells like proposed by Siemon et al. [26]. Given the fixed locations of these grid cells, and potentially overlapping bounding boxes of detected objects, the primary objective of the discrete BN is to analyze dependencies between the static cells and dynamic objects in terms of conditional probability distributions. This is accomplished by modeling the dependency relationship via several characteristics of the bounding box. This subsection describes those characteristics, the discretization proposal of some, and is followed by the design of the resulting network as such. Before that, however, it is important to clarify some of the relevant terminology.
Bayes’ Theorem is a means to mathematically determine the conditional probability of an event based on some prior knowledge, which is relevant to that event. These events are also commonly called observations, and refer to attributes which span the sample space $\mathcal{S}$ of the problem. Each such attribute is further called random variable (RV) and has a fixed set of possible values, known as its value space (VS) which may be modeled by either discrete or continuous values. This section enumerates and describes each of the RVs in $\mathcal{S}$, their respective types (numerical or categorical) and VSs considered relevant to the task at hand. This lays the groundwork for the subsequent description of their interdependencies. We use domain knowledge from the field of image processing to further reason over the joint probability distribution spanned by our problem domain. More specifically, we impose a structure of interdependencies between the high-level appearance/motion attributes of a bounding box and its location within a frame.
@@ 3.1.1 Random Variables for Anomaly Detection in Videos
In the description of all defined RVs and their corresponding VSs that follows in Table 1 it can be seen that some VSs were defined in the numeric space while the rest is categorical. The motivation behind this modelling choice was purely to provide a high-level understanding of what the values of all these RVs statistically represent. For instance, we claim that it is much easier for a human to understand the conceptual difference of different bounding box sizes when comparing a small box to a large one. Even though the same applies to all object classes that are considered in this work, we refrain from listing all $80$ categories of the MS-COCO dataset in the list below and use their individual numerical identifiers instead.
@@
All categorical VSs are determined statistically by the standard deviation $\sigma$ and the mean $\mu$ of the underlying class-wise distribution of the given RV as observed during training. The data is discretized based on the observation’s distance from the mean, expressed through multiples of $\sigma$. Regarding the temporal RVs defined in Table 1, it is left to mention that they are determined by the displacement between the two geometric center coordinates of the corresponding bounding boxes as the object is tracked across two consecutive frames. A visualization of the set of all spatial RVs is provided in Figure 2(a) for better understanding.
@@ 3.1.2 Graphical Representation
A well-informed choice of nodes and proper modeling of the dependency structures within the discrete BN are crucial for the model to be conditioned on prior knowledge. Such knowledge may depend on a variety of contexts relevant to dataset, scene, or maybe recording perspective of the video footage. In the following, we aim to contrast a purely spatial and a spatio-temporal network version. In this way, we underline the extension capabilities of the discrete BN from capturing purely appearance-based anomalies towards a model that is also capable of detecting motion-based outliers in a video stream. Figure 2 thus illustrates two different network versions, modeled for this purpose in the form of Directed Acyclic Graphs (DAGs) with unidirectional dependency connections between nodes such that two nodes are connected by at most one edge. Throughout the remainder of this paper they will be referred to as: spatial and spatio-temporal network models.
@@
(b) Spatial Model
(c) Spatio-Temporal Model
Figure 2: From left to right: All spatial RVs from Table 1 are illustrated on a sample image from StreetScene which was converted to greyscale for better visualization purposes. Different network versions to model conditional relations between all RVs in spatial and spatio-temporal task settings are contrasted.
As it can be seen in Figure 2(b) and Figure 2(c), both network constellations share the same spatial dimension including the corresponding nodes. Here, modelling the dependency structure as given by the directed edges is motivated as follows: In order to give the problem domain we aim to represent a global context the frame node F is serving the purpose of the root node in the graph. While each frame is divided into a uniform grid structure composed of grid cells G, the intersection I between an object’s bounding box and the respective grid cell G is constrained by the aspect ratio BAR and the size of the box BS. The latter in combination with drawing a dependency between grid cell G and the bounding box size BS aims at encoding the scene depth given the influence it can have on the perceived size of some observed object. Following this modelling approach, both, the size BS and aspect ratio BAR of a bounding box are constrained according to the class the observed object is representing. To illustrate this with an example, it can be claimed that the bounding box sizes of vehicles are on average larger than those drawn around humans, including the fact the the most common aspect ratio of a car resembles on average the landscape format. The bounding box of a human captured on CCTV footage in the city on the other hand would most likely be assumed to be of portrait format.
Compared to Figure 2(b), however, the network shown in Figure 2(c) consists of two additional RVs which span the temporal dimension of the model, velocity V and direction D of the object. Following the notion of encoding scene depth in our model, the velocity V of a bounding box is not only constrained on the object class (humans tend to move slower than cars, e.g.) but also on the area of the frame in which it moving (planes flying in the sky appear to move relatively slow when observed from the ground, e.g.). Proof for how the addition of a temporal dimension to the discrete BN in contrast to its spatial counterpart can effectively enhance the performance of the model in terms of detecting spatio-temporal anomalies is given further below in Figure 3.
@@ 4.1.1 Frames
Results presented in Table 2 are based on experiments that involved either every third training frame (CUHK Avenue) or every fifth one (ShanghaiTech and StreetScene). Reducing the amount of training images was primarily motivated by statistical models being prone to over-fitting when faced with too much training data. Additionally, it naturally implies less processing time during both, training and inference. Both aspects are supported by corresponding data presented in the supplementary material in Section 6.
@@ 4.2 Data Preprocessing
We utilize BoT-SORT [2] to pre-process all training and test frames. We choose it to be based on the object detector YOLOv7, pretrained on the MS-COCO dataset, including an object re-identification module that was trained on the Multiple Object Tracking 17 (MOT17) dataset [19]. In order to alter the tracker’s capabilities of catching very small objects for the high angle shot dataset StreetScene we additionally finetune the object detector on the VisDrone-DET[30] dataset which belongs to the collection of datasets recorded out of Unmanned Aerial Vehicles (UAV) such as drones from extreme bird’s eye view perspectives.
@@ 4.2.1 Confidence Threshold
Avoiding unnecessary noise in the discrete BN model, we filtered all objects coming from the MOT module according their corresponding confidence score given a certain threshold. This threshold was determined in accordance with the overall distribution of confidence scores of boxes detected during training on a particular scene. It was set dynamically prior to the generation of observations: one for the class person, and another one for all remaining classes. This differentiation was motivated by the fact that common object detectors are strongly biased towards detecting human beings. Various experiments showed that the confidence threshold yielding best results across all datasets is approximately $2\cdot\sigma$ smaller than its mean.
@@ 4.3 Training
By using the term training, we refer to the estimation of the joint probability distribution spanned by $\mathcal{S}$ described in Table 1. The data used for this purpose is fully discrete, and can therefore be encoded in a table of observations. Following the design of the network structures presented in Figure 2, a single observation consists of a fixed set of features comprising assignments to the RVs of $\mathcal{S}$. For any grid cell intersecting with any bounding box in a given frame, one such observation is created and added to our tabular dataset. We process the bottom part of a bounding box which affects only a proper subset of cells in a particular row of the grid structure. This is motivated by the finding that a full-box approach can result in excessive noise when confronted with significant occlusions between objects. We supply corresponding proof in the supplementary material in Section 6.
@@ 4.4 Technical Details
In its current implementation the training and prediction phase of our PGMs is conducted on the CPU mentioned above. Reaching a processing speed of almost up to 30 cells per second during inference while the training completes in less than a second (both are dataset dependent), we refer to the supplementary material in Section 6 for further details on this note.
4.5 Evaluation
4.5.1 Metrics
In choosing evaluation metrics, we follow the majority of published research on VAD: Frame-level Area Under the Curve (AUC), and Region-Based Detection Criterion (RBDC) and Track-Based Detection Criterion (TBDC). The latter two have been proposed by Ramachandra and Jones [22] in order to replace the previously used frame-based criterion because it is neither capable of locating an anomaly within a frame nor accounting for different amounts of anomalous regions. Similarly to Singh et al. [27], we report frame-level AUC scores nevertheless, for completeness reasons. The calculation of both RBDC and TBDC scores has been facilitated by a script provided by Georgescu et al. alongside their work described [11]. Given the large discrepancy between RBDC and TBDC scores and to thus allow for a simpler comparison of the said metrics, we add an additional column to our results tables containing their average value.
@@ 4.6 Quantitative Results
Results in Table 2 we achieved on all benchmark datasets indicate highly competitive performance of our proposed approach on CUHK Avenue which is solely based on the analysis of high-level bounding box features and their particular location with the frame. Our model achieves $66.14\%$ on RBDC/TDBC average which marks the third-best result to date. In Figure 3, we also illustrate the effect of transitioning from the spatial model shown in Figure 2(b) to the spatio-temporal version in Figure 2(c). We chose test videos #03, 04 and 07 of CUHK Avenue for this demonstration purpose because of the little background noise and the clearly visible temporal anomalies they contain. The enhanced capability of detecting temporal anomalies with our spatio-temporal model version is indicated here by the clear drop in frame-level AUC in areas highlighted with red background, i.e., anomalous frames.
Table 2: Accuracy Scores (%): Frame-level AUC, RBDC, TBDC, and the mean $\mu_{\text{R,T}}$ of the two latter. We report accuracies of model Figure 2(c). Stressing highest, second-highest and third-highest results per metric.
@@
Similarly to the previous benchmark dataset, the results achieved on ShanghaiTech remain competitive. Scoring third-best in terms of RBDC scores reported to date on this dataset also makes this method a close runner-up on RBDC/TBDC average. Further, despite the level of difficulty introduced by StreetScene, we set a new SOTA baseline with our spatio-temporal discrete BN model, and report $30.65\%$ in RBDC and $66.03\%$ in TBDC, surpassing Singh et al. [27] by nearly $4\%$ on RBDC/TBDC average.
Figure 3: Contrasting spatial (Figure 2(b)) and spatio-temporal model versions (Figure 2(c)) based on three concatenated CUHK Avenue test videos, containing only temporal anomalies (a man running, a child jumping) in terms of frame-level AUC scores. The images are generated by the spatio-temporal model version.
@@ 4.7 Qualitative Results: Explainability
Given an object of very low probability, i.e., an anomaly, observed in some cell during test time, we aim at finding a statistical explanation for it. To this end we present an example visualization of an anomaly taken from the third test video of CUHK Avenue in Figure 4. Based on the breakdown of individual probability distributions of all visual attributes we defined in Table 1, it can be deducted from these charts that apart from the velocity direction ("SW"), it is primarily the size ("x-large") and velocity ("lightning fast") of the given bounding box which are contributing most to the object’s low probability. The reason for this is that first, the measured values within those two categories do not correspond to the most likely ones, i.e., the ones occupying the largest fraction of the respective chart, and that second, they in fact represent the least likely values among all others that were encountered during training. Every chart was generated based on the output of a query which was supplied with cell index and all bounding box attributes except for the one under investigation as hard evidence. We provide equal visualizations for the remaining two cells that the bounding box overlaps with in the supplementary material in Section 6.
Figure 4: A visualization proposal of explaining the anomaly score extracted for a specific cell given the appearance of an intersecting bounding box using the RVs defined in Table 1.
@@ 5 Conclusions and Future Work
In conclusion, throughout the experiments conducted, PGMs proved to be very suitable candidates to tackle VAD when combined with most recent MOT models, offering a wide range of graph modelling opportunities. While many other works are challenged by StreetScene, we have shown that the modeling framework we used is sufficiently capable to cope with the scene it depicts delivering new SOTA baselines on this dataset. Additionally, with the competitive performance achieved on CUHK Avenue and ShanghaiTech we have shown that the representation of objects via their surrounding bounding boxes is indeed sufficient to successfully detect anomalies in videos. Another advantage of deployed ML models is that they are capable of performing better with less training data compared to the entire training set which we also conclude based on our experiments. We also foresee that, through these means, relative advantages can be found in computational complexity, and both space and time demands.
@@
The first aspect we would like to address concerns the fact that the deployed PGM is operating on discretized variable values. While this can be embedded in a relatively stable statistical setup, it has limitations when compared to a purely continuous dimension. The exploration of continuous PGMs for VAD is thus left for future work. Second, given the rigid grid structure which is imposed on the video frames, it is debatable whether there might exist more efficient alternatives to induce localization capabilities to the model, and is therefore also left for future work. Third, as far as our experiments are concerned, the pipeline is only as good as the closed-set tracking module used, with its error rate being propagated to the PGM. We leave the validation of this hypothesis for future work, on the assumption that a VAD dataset with GT tracking labels for training comes to be available. This includes the implementation of open-world object tracking and/or detection modules.

### v2

@@ Abstract
In this study, we formulate the task of Video Anomaly Detection as a probabilistic analysis of object bounding boxes. We hypothesize that the representation of objects via their bounding boxes only, can be sufficient to successfully identify anomalous events in a scene. The implied value of this approach is increased object anonymization, faster model training and fewer computational resources. This can particularly benefit applications within video surveillance running on edge devices such as cameras. We design our model based on human reasoning which lends itself to explaining model output in human-understandable terms. Meanwhile, the slowest model trains within less than $7$ seconds on a $11^{th}$ Generation Intel Core i9 Processor. While our approach constitutes a drastic reduction of problem feature space in comparison with prior art, we show that this does not result in a reduction in performance: the results we report are highly competitive on the benchmark datasets CUHK Avenue and ShanghaiTech, and significantly exceed on the latest State-of-the-Art results on StreetScene, which has so far proven to be the most challenging VAD dataset. We release our code to the community at: https://github.com/milestonesys-research/VAD-with-PGMs/.
@@ 1 Introduction
  • A new, simplified perspective on problem domain representation within VAD.
  • A novel approach for fully object-centric VAD in which 8 high-level object bounding box attributes are learned by a discrete Bayesian Network to detect visual outliers in video streams.
  • Human-understandable model and result explainability within the context of the high-level attributes used to model the problem domain.
@@ 3.1 Discrete Bayesian Network
The foundation of our study is a discrete BN which facilitates probabilistic reasoning over observations made in our VAD-specific problem domain. Each such observation is composed of a set of different variables that can independently influence the final outcome of this process. Embedded in Bayes’ Theorem, each such variable is called random variable (RV) and has a fixed set of possible values, known as its value space (VS). Together, all RVs span the sample space $\mathcal{S}$. The calculation of the resulting joint probability distribution to detect outliers is governed by domain-specific dependency structures. A well-informed choice of these variables and proper modeling of the dependency structures are crucial for any BN to be successfully conditioned on prior knowledge. Within VAD, such knowledge may depend on a variety of contexts relevant to dataset, scene, or maybe even recording perspective of the video footage. In this approach, we encode such information in our model via $8$ spatio-temporal variables which are derived from high-level appearance/motion attributes of a bounding box and its location within a frame.
@@ 3.1.1 Random Variables for Anomaly Detection in Videos
In the description of all defined RVs and their corresponding VSs that follows in Table 1 it can be seen that some VSs were defined in the numeric space while the rest is categorical. The motivation behind this modelling choice was purely to provide a high-level understanding of what the values of all these RVs statistically represent. For instance, we claim that it is much easier for a human to understand the conceptual difference between a small and a large bounding box. Even though the same applies to all object classes that are considered in this work, we refrain from listing all $80$ categories of the MS-COCO dataset in the list below and use their individual numerical identifiers instead. To further provide the BN with the capability of localizing anomalous events within a single frame, we divide the image into a uniform grid structure of quadratic cells like proposed by Siemon et al. [26].
@@
The thresholds for all categorical VSs are determined statistically per dataset by the standard deviation $\sigma$ and the mean $\mu$ of the underlying class-wise distribution of the given RV as observed during training. The data is discretized based on the observation’s distance from the mean, expressed through multiples of $\sigma$. Regarding the temporal RVs defined in Table 1, it is left to mention that they are determined by the displacement between the two geometric center coordinates of the corresponding bounding boxes as the object is tracked across two consecutive frames. A visualization of the set of all spatial RVs is provided in Figure 2(a) for better understanding.
@@ 3.1.2 Graphical Representation
We now provide insight into how we express VAD-specific domain knowledge in terms of interdependencies between all RVs given in Table 1. This knowledge entails spatial and temporal clues to address appearance- and motion-based anomalies in videos. The resulting BN is given in the form of a Directed Acyclic Graph (DAGs) with unidirectional dependency connections between nodes (representative for RVs) such that two nodes are connected by at most one edge. Modelling the dependency structure as given by the directed edges in Figure 2(b) is motivated as follows: In order to give the problem domain we aim to represent a global context the frame node F is serving the purpose of the root node in the graph. While each frame is divided into a uniform grid structure composed of grid cells G, the intersection I between an object’s bounding box and the respective grid cell G is constrained by the aspect ratio BAR and the size of the box BS. Drawing a dependency between grid cell G and the bounding box size BS aims at encoding the scene depth given the influence it can have on the perceived size of some observed object. Following this modelling approach, both the size BS and aspect ratio BAR of a bounding box are constrained according to the class the observed object is representing. To illustrate this with an example, it can be claimed that the bounding box sizes of vehicles are on average larger than those drawn around humans, including the fact the the most common aspect ratio of a car resembles on average the landscape format. The bounding box of a human captured on CCTV footage in the city on the other hand would most likely be assumed to be of portrait format.
We express the temporal dimension of the model via the velocity V and direction D of the object. Following the notion of encoding scene depth in our model, the velocity V of a bounding box is not only constrained on the object class (humans tend to move slower than cars, e.g.) but also on the area of the frame in which it is moving (planes flying in the sky appear to move relatively slow when observed from the ground, e.g.).
@@
(b) Spatio-Temporal Model
Figure 2: Left: All spatial RVs from Table 1 except for the frame F are illustrated on a sample image from StreetScene which was converted to greyscale for better visualization purposes. Right: Our proposed BN model with conditional relations between all RVs to perform VAD.
@@ 4.1.1 Frames
Results presented in Table 2 are based on experiments that involved either every third training frame (CUHK Avenue) or every fifth one (ShanghaiTech and StreetScene). Reducing the amount of training images was primarily motivated by statistical models being prone to over-fitting when faced with too much training data. Additionally, it naturally implies less processing time during both, training and inference. Both aspects are supported by corresponding data presented in the supplementary material.
@@ 4.2 Data Preprocessing
We utilize BoT-SORT to pre-process all training and test frames. We choose it to be based on the object detector YOLOv7, pretrained on the MS-COCO dataset, including an object re-identification module that was trained on the Multiple Object Tracking 17 (MOT17) dataset [19]. In order to alter the tracker’s capabilities of catching very small objects in video frames we additionally finetune the object detector on the VisDrone-DET[30] dataset which belongs to the collection of datasets recorded out of Unmanned Aerial Vehicles (UAV) such as drones from extreme bird’s eye view perspectives.
@@ 4.2.1 Confidence Threshold
Avoiding unnecessary noise in the discrete BN model, we filtered all objects coming from the MOT module according their corresponding confidence score given a certain threshold. This threshold was determined in accordance with the overall distribution of confidence scores of boxes detected during training on a particular scene and per dataset. It was set dynamically prior to the generation of observations: one for the class person, and another one for all remaining classes. This differentiation was motivated by the fact that common object detectors are strongly biased towards detecting human beings. Various experiments showed that the confidence threshold yielding best results across all datasets is approximately $2\cdot\sigma$ smaller than its mean $\mu$.
@@ 4.3 Training
By using the term training, we refer to the estimation of the joint probability distribution spanned by $\mathcal{S}$ described in Table 1. The data used for this purpose is fully discrete, and can therefore be encoded in a table of observations. Following the design of the network structure presented in Figure 2(b), a single observation consists of a fixed set of features comprising assignments to the RVs of $\mathcal{S}$. For any grid cell intersecting with any bounding box in a given frame, one such observation is created and added to our tabular dataset. We process the bottom part of a bounding box which affects only a proper subset of cells in a particular row of the grid structure. This is motivated by the finding that a full-box approach can result in excessive noise when confronted with significant occlusions between objects. We supply corresponding proof in the supplementary material.
@@ 4.4 Technical Details
In its current implementation the training and prediction phase of our PGMs is conducted on the CPU mentioned above. Reaching a processing speed of almost up to 30 cells per second during inference while the training completes in less than a second (both are dataset dependent), we refer to the supplementary material for further details on this note.
4.5 Evaluation Metrics
In choosing evaluation metrics, we follow the majority of published research on VAD: Frame-level Area Under the Curve (AUC), and Region-Based Detection Criterion (RBDC) and Track-Based Detection Criterion (TBDC). The latter two have been proposed by Ramachandra and Jones [22] in order to replace the previously used frame-based criterion because it is neither capable of locating an anomaly within a frame nor accounting for different amounts of anomalous regions. This metric is thus only roughly indicative of a method’s true performance. Similarly to Singh et al. [27], we report frame-level AUC scores nevertheless, for completeness reasons. The calculation of both RBDC and TBDC scores has been facilitated by a script provided by Georgescu et al. alongside their work described in [11]. Given the large discrepancy between RBDC and TBDC scores and to thus allow for a simpler comparison of the said metrics, we add an additional column to Table 2 containing their average value.
@@ 4.6 Quantitative Results
Results shown in Table 2 indicate strong performance of our proposed approach across all three benchmark datasets especially given that it is solely based on the analysis of high-level bounding box features and their particular location with the frame. Our model achieves $66.14\%$ on RBDC/TDBC average which marks the third-best result to date on CUHK Avenue. Similarly, the results achieved on ShanghaiTech remain competitive. Scoring third-best in terms of RBDC scores reported to date on this dataset also makes this method a close runner-up on RBDC/TBDC average. Further, despite the level of difficulty introduced by StreetScene, we set a new SOTA baseline with our spatio-temporal discrete BN model, and report $30.65\%$ in RBDC and $66.03\%$ in TBDC, surpassing Singh et al. [27] by nearly $4\%$ on RBDC/TBDC average.
Table 2: Accuracy Scores (%): Frame-level AUC, RBDC, TBDC, and the mean $\mu_{\text{R,T}}$ of the two latter. Stressing highest, second-highest and third-highest results per metric.
@@ 4.7 Qualitative Results: Explainability
Given an object of very low probability, i.e., an anomaly, observed in some cells during test time, we aim at finding a statistical explanation for it. To this end, we present an example visualization of an anomaly taken from test video $11$ of the benchmark dataset ShanghaiTech in Figure 3. Based on the breakdown of individual probability distributions of all visual attributes we defined in Table 1, it can be deducted that the detected object (’bike’) was flagged as anomalous in the given area due to the observed values of attributes Class (’bicycle’), BoxSize (’xlarge’), Velocity (’lightning fast’) and Direction (’E’). More likely values for these categories based on the training data, highlighted with a thick orange border, would correspond to ’bench’, ’xsmall’, ’normal’ and ’SW’, respectively. In other words, the object is more likely to be a ’bench’ under the remaining observed evidence. If it indeed is a ’bicycle’, however, then the movement direction is more likely to be ’SW’ given the remaining evidence, and so on. The charts show the average values of extracted probabilities per cell intersecting with the bottom border of the object bounding box. The total corresponding frame area is highlighted with a pink rectangle. We provide similar explainability visualizations for test samples from the remaining benchmark datasets in the supplementary material.
Figure 3: A visualization proposal of explaining the anomaly score extracted for a set of cells given the current appearance/velocity expressed of the bounding box through RVs defined in Table 1.
4.8 Ablation Study
In the following, we want to emphasize the effect of the temporal dimension in our discrete BN shown in Figure 2(b) on the detection of motion-based anomalies in videos. To this end, we create a purely spatial counterpart by removing the temporal dimension of our spatio-temporal model only keeping the inner spatial dimension. To contrast the performances of both network versions, we generated Figure 4 based on three concatenated test videos. We chose test videos #03, 04 and 07 of CUHK Avenue for this demonstration purpose because of the little background noise and the clearly visible temporal anomalies they contain. The enhanced capability of detecting temporal anomalies with our spatio-temporal model version is indicated here by the clear drop in frame-level AUC in areas highlighted with red background, i.e., anomalous frames.
Figure 4: Contrasting spatial and spatio-temporal model versions (Figure 2(b)) based on three concatenated CUHK Avenue test videos, containing only temporal anomalies (a man running, a child jumping) in terms of frame-level AUC scores. The images are generated by the spatio-temporal model version.
@@ 5 Conclusions and Future Work
In conclusion, PGMs proved to be very suitable candidates to tackle VAD when combined with most recent MOT models. While many other works are challenged by StreetScene, we have shown that the modeling framework we used delivers new SOTA baselines on all recognized metrics on this dataset. We hypothesize that our model is not capable of beating previous SOTA results on CUHK Avenue and ShanghaiTech as opposed to StreetScene because of the anomaly types the former two datasets contain. Often, these anomalies depict anomalous human behavior such as throwing papers in CUHK Avenue, and dropping/playing with objects in ShanghaiTech. Unless such behavior results in a significant change of observed RV values, our method will not be capable of detecting such anomalies. Nevertheless, we believe that the results on all benchmark datasets prove that the representation of objects via their surrounding bounding boxes is indeed sufficient to successfully detect anomalies in videos, representing a novel and valuable insight into the domain of VAD. Another advantage of deployed ML models is that they are capable of performing better with less training data which we also conclude based on our experiments. Through these means, we foresee that relative advantages can be found in computational complexity, and both space and time demands.
@@
The first aspect we would like to address concerns the fact that the deployed PGM is operating on discretized variable values. While this can be embedded in a relatively stable statistical setup, it has limitations when compared to a purely continuous dimension. The exploration of continuous PGMs for VAD is thus left for future work. Second, given the rigid grid structure which is imposed on the video frames, it is debatable whether there might exist more efficient alternatives to induce localization capabilities to the model, and is therefore also left for future work. Third, as far as our experiments are concerned, the pipeline is only as good as the closed-set tracking module used, with its error rate being propagated to the PGM. We leave the validation of this hypothesis for future work, on the assumption that a single-scene VAD dataset with GT tracking labels for training comes to be available. This includes the implementation of open-world object tracking and/or detection modules.